Not sure if I am posting this in the correct area so please accept my apologies in advance, im struggling with the following:
Transpose the following formula to make x the subject:
y= 7+x/14
I think it works something like this to move the 7 accross i takeaway 7 from both sides and the 14 i multiply on both sides
so i come up with x=14y-7
Is this correct and can you suggest where i can find further reading on this subject?

Answer by CubeyThePenguin(3113)

y = 7 + x/14
y - 7 = x/14
14(y - 7) = x
14y - 126 = x

You just needed to distribute the 14 in 14(y - 7). Every other step is good.